# Evaluating Decision Optimality of Autonomous Driving via Metamorphic Testing

### v1

@@ 5.1. RQ1: Effectiveness of Decictor
In summary, the comparison of Decictor with Random and Random-$\delta$ demonstrates its effectiveness in detecting NoDSs. When contrasted with the three safety-oriented baselines, the results emphasize the importance of robust decision-making testing. In fact, Decictor and existing safety-oriented baselines are complementary, given their distinct testing objectives.
Table 2. Comparison of mutations.
Table 3. Comparison of feedback.
Method %Mutationnormal-↑\,$\uparrow$ #NoDSnormal-↑\,$\uparrow$
S1 S2 S3 S4 Avg. S1 S2 S3 S4 Sum
w/o Cons 71.7 71.3 74.8 65.5 70.8 2.7 4.5 1.6 5.7 14.5
w/o Mot 71.1 80.9 98.6 68.9 79.9 4.0 8.6 5.2 10.8 28.6
w/o Rem 97.5 97.3 99.4 97.2 97.9 3.9 9.3 4.9 9.1 27.2
Decictor 98.0 97.5 99.4 97.5 98.1 9.4 11.9 9.0 15.7 46.0
Method #NoDS
S1 S2 S3 S4 Sum
F-Random 3.0 5.1 4.4 7.2 19.7
F-MR 3.0 6.7 1.2 7.5 18.4
F-Path 4.3 10.8 5.0 10.6 30.7
F-Behavior 5.3 9.0 5.4 11.5 31.2
Decictor 9.4 11.9 9.0 15.7 46.0
Table 3. Comparison of feedback.
5.1.2. Case Study.
Figure 6 showcases four NoDS examples generated from the four initial ODSs. Each example includes three scenarios: the initial ODS, the identified NoDS, and the reproduced scenario (i.e., the optimal decision in the NoDS). The examples illustrate that the ego vehicle takes non-optimal paths in the detected NoDSs, even though the optimal ones are still available. Details are explained as follows:
@@
Case 1. Inaccurate Prediction on Vehicle Status. Decictor adds a new NPC vehicle with a low initial speed in the NoDS. Despite being initially placed on the ego vehicle’s optimal path and far away, the NPC vehicle will have moved by the time the ego vehicle reaches that segment. The original optimal path remains available, as shown in the reproduced scenario. However, the ADS makes an incorrect prediction, considering the NPC vehicle as stationary. Consequently, it stops lane changing and chooses a less efficient path.
@@
Case 2. Inaccurate Prediction on Obstacle Impact. Decictor introduces two obstacles along the lane boundaries near the exit of the junction. These two obstacles do not influence the optimal path. In this NoDS, the ego vehicle incorrectly assesses that the target lane is impassable. Consequently, it chooses to approach the destination from an adjacent, external lane. This choice is not optimal and introduces significant risks within the context of real-world traffic.
Case 3. Inaccurate Prediction on Safe Regions. Decictor introduces five static obstacles in the NoDS, without disrupting the optimal path (as shown in the reproduced scenario). However, the ego vehicle deviates from the optimal path and chooses an alternative path to reach the destination. It involves navigating around the obstacles and changing lanes four times, leading to a non-smooth and dangerous motion. This scenario illustrates the ADS’s insufficient intelligence, resulting in non-optimal decisions.
Case 4. Inaccurate Prediction on Vehicle Intention. Decictor introduces a new NPC vehicle in the NoDS. Note that this NPC vehicle does not intersect with the optimal path of the ego vehicle. The NPC vehicle travels along the blue path at a low speed. However, the ego vehicle incorrectly predicts that the NPC vehicle intends to change to the left lane. This misinterpretation causes the ego vehicle to focus excessively on the NPC’s movement, leading it to overlook the inadequate available space to execute a left turn. As a result, the ego vehicle is compelled to select a non-optimal path to reach its destination.
Answer to RQ1: Decictor significantly outperforms existing methods in identifying NoDSs. While conventional ADS testing methods primarily focus on detecting safety issues, Decictor can serve as a complement to them for a more comprehensive evaluation of ADSs.
5.2. RQ2: Usefulness of Mutation and Feedback
@@
We assess the usefulness of the key components in Decictor, i.e., the non-invasive mutation and the feedback mechanism. To achieve this, we conducted a thorough evaluation by configuring a series of variants of Decictor and then proceeded to evaluate their usefulness.
5.2.1. Mutation.
@@
For mutation, we compared Decictor with its three variants: (1) w/o Cons applies a random mutation instead of the non-invasive mutation in Decictor, aiming to evaluate the effectiveness of the non-invasive mutation; (2) w/o Mot sets the time step to 0s in the calculation of non-invasive feasible areas (i.e., only considering the motion constraint at each timestamp), aiming to measure the influence of the motion constraint between two successive timestamps. (3) w/o Rem uses only the adding operation in Decictor under the constraint of non-invasive mutation, aiming to assess the effectiveness of the combination of the two mutation operations, i.e., adding and removing participants. As shown in Table 3, we find that w/o Cons generates the fewest valid mutations (70.8%) and detects the smallest number of NoDSs (14.5), underlying the importance of the non-invasive mutation. Comparing w/o Mot with Decictor, we observed that only considering the motion constraint at each timestamp is insufficient, and the continuous-time motion constraint plays a significant role in calculating the non-invasive feasible area. Our preliminary experiments revealed that %Mutation increases while the #NoDS decreases as the time step increases. The reason is that a long time step will yield scenarios with reduced interactivity between the ego vehicle and the added participant, thus exerting a lower impact on the decision-making process of the ADS. Comparing w/o Rem and Decictor, we found that mutation with only the adding operation is more likely to induce more invalid mutations and a lower number of NoDSs. This is because the adding operation brings too many obstacles and may cause the failure of the motion task. Thus, the removing operation is also important for generating NoDSs.
5.2.2. Feedback.
@@
For feedback, we implemented four variants: 1) F-Random replaces the feedback of Decictor with a random selection (from $\textbf{Q}\cup\textbf{Q}^{\prime}$) to evaluate the effectiveness of our feedback strategy. 2) F-MR replaces the feedback of Decictor with the grid similarity used in our MR checking (see Equation 6) to compare the performance between using a coarse-grained grid distance and a fine-grained path distance. 3) F-Path and F-Behavior consider only the driving path feedback and the behavior feedback, respectively, to evaluate the usefulness of either feedback type.
@@
Table 3 shows the experimental results of #NoDSs. The comparative results between F-Random and Decictor (19.7 vs. 46.0 in total) illustrate the effectiveness of the feedback used in Decictor. The results of F-MR (18.4) indicates that the grid similarity is not an effective feedback metric, as it may overlook some scenarios with a high probability of inducing NoDSs. From the ablation results of F-Path and F-Behavior (30.7 and 31.2, respectively), we see that both behavior feedback and driving path feedback are beneficial for detecting NoDSs. Their combination achieves the best performance.
@@
5.3. RQ3: Test Efficiency of Decictor
Figure 7 shows the cumulative number of NoDSs over the execution time of different methods. We can find that as Decictor continues its execution, the detection of NoDSs steadily increases, whereas other methods quickly reach a stable state. We further assess the time performance of different components in Decictor, including the overhead of mutation, oracle checking, feedback calculation, and simulation. Specifically, we analyze the average time to handle a scenario. The results are summarized in Table 4, where the columns Mutation, Oracle Check, Feedback, and Simulation represent the average time taken by a scenario for mutation, validity checking, fitness computation, and running the scenario in the simulation platform, respectively. 0.01* represents that the time cost is small, i.e., less than 0.01. For Random, the oracle and feedback stages are not involved, being marked as ‘N/A’.
The overall results show that, for all tools, the simulation process spends the majority of time. For instance, on average, Decictor takes 77.58 seconds to process a scenario, where 67.20 seconds (86.62%) are used in running the scenario. Decictor spends slightly more time in Mutation and Oracle Check than others due to the computation of the non-invasive feasible area for each waypoint and the grid-based similarity-checking mechanism. However, the computation time of Decictor remains within an acceptable range. SAMOTA consumes the most time in the feedback due to the necessity of training a surrogate model. The differing time durations for simulation across various tools can be attributed to the different scenarios each tool generates (e.g., time-out scenarios), which in turn require varying amounts of simulation time.
Table 4. Results of time performance (s)
@@ 5.3. RQ3: Test Efficiency of Decictor
Figure 7. #NoDSs over the execution of different methods.
5.4. Threats to Validity
@@
Decictor suffers from some threats. The selection of the initial ODSs could potentially influence the results. To mitigate this, we have judiciously selected four motion tasks and created corresponding scenarios. Each scenario was run multiple times, with the optimal motion validated through both numerical analysis and expert judgment. Despite the MR Check $\epsilon$ in Decictor, the manual confirmation of NoDS-Hum may present another threat as different people may identify a scenario as either an ODS or a NoDS. To counter this, on one hand, we carefully select the threshold $\varepsilon$ in the MR Check and $\Delta t$ in the computation of non-invasive feasible area to reduce the number of such scenarios; on the other hand, to ensure robustness and consistency, all co-authors independently review the generated NoDSs, and only those confirmed by unanimous agreement among all authors will be considered in NoDS-Hum. Randomness may also be a threat to our results, and we mitigate the threat by repeating the experiment multiple times. Another potential threat could be that our current MR primarily addresses non-optimal decisions related to selecting inefficient paths. However, there can be other types of non-optimal decisions, such as lack of smooth driving or unnecessary jittery movements, which are not considered by our current MR. In the future, we plan to extend the MR (e.g., considering velocity and acceleration) to detect more non-optimal decisions. Lastly, the customization of baselines to the Apollo+SimControl simulation environment could potentially affect the results. However, we only modified the interface between the algorithms and the simulation environment, leaving the algorithms themselves untouched. Moreover, to ensure the correctness, all co-authors thoroughly reviewed the code.
6. Related Work
@@
Safety-Guided ADS Testing. Safety plays the most important role in the development of ADSs. However, guaranteeing everlasting safety is challenging and many approaches have been proposed to generate safety-critical scenarios for evaluating the safety of ADSs. They can be divided into data-driven approaches (Zhang and Cai, 2023; Deng et al., 2022b; Gambi et al., 2019a; Najm et al., 2013; Nitsche et al., 2017; Roesener et al., 2016; Paardekooper et al., 2019) and searching-based approaches (Cheng et al., 2023; Hildebrandt et al., 2023; Gambi et al., 2019b; Han et al., 2021; Li et al., 2020; Haq et al., 2022; Zhong et al., 2023; Tang et al., 2021c; Zhou et al., 2023; Tang et al., 2021b, a). Data-driving methods produce critical scenarios from real-world data, such as traffic recordings (Deng et al., 2022b; Paardekooper et al., 2019; Roesener et al., 2016) and accident reports (Gambi et al., 2019a; Najm et al., 2013; Nitsche et al., 2017; Zhang and Cai, 2023). Search-based methods use various technologies to search for safety-critical scenarios from the scenario space, such as guided fuzzing (Cheng et al., 2023; Li et al., 2020; Pang et al., 2022), evolutionary algorithms (Gambi et al., 2019b; Han et al., 2021; Tang et al., 2021a, b, c; Zhou et al., 2023), surrogate models (Haq et al., 2022; Zhong et al., 2023), reinforcement learning (Haq et al., 2023; Feng et al., 2023) and reachability analysis (Hildebrandt et al., 2023; Althoff and Lutz, 2018). All these methods aim to generate critical scenarios for evaluating ADSs’ safety-critical requirements, such as collision avoidance and reaching the destination. In contrast, our work focuses on evaluating the non-safety-critical requirements of ADSs, which is also crucial for the real-world deployment of ADSs. To the best of our knowledge, it is the first one considering the non-safety-critical requirements of ADSs.
Metamorphic Testing for ADS. Metamorphic testing (MT) has been an emerging approach to generating test cases and overcoming the oracle problem (Chen et al., 2018; Xie et al., 2011). Recently, MT has also been applied to test deep learning systems, such as the perception system of ADSs (Xie et al., 2022) and end-to-end deep learning-based ADSs (Zhang et al., 2018; Wang and Su, 2021; Deng et al., 2022a). For example, the authors in (Wang and Su, 2021) designed an MR to test the robustness of an object detection system, which can be roughly described as: Inserting an additional object that does not overlap with preexisting objects in the source image should not affect the detection of the preexisting objects in the new image. However, there is only a little work focusing on MT for multi-module ADSs (Zhou and Sun, 2019; Han and Zhou, 2020). In (Han and Zhou, 2020), the authors propose an MT method to identify avoidable collision scenarios, which means there are genuine failures in the ADS under test. In contrast to existing MT techniques that primarily focus on the correctness and safety of the system under test, we propose an MT method to detect vital non-safety-critical violations of the ADS under test.
7. Conclusion

### v3

@@ V-A1 Comparative Results
In summary, the comparison of Decictor with Random and Random-$\delta$ demonstrates its effectiveness in detecting NoDSs. When contrasted with the seven state-of-the-art baselines, the results emphasize the importance of robust path-planning decision testing. Decictor and these safety-oriented baselines are complementary, given their distinct testing objectives.
Figure 6: Cases of NoDSs detected by Decictor. The red vehicle is the AV. Optimal paths are indicated Solid lines, while non-optimal paths are displayed in Dotted lines. The first column is the initial ODSs, the second column shows the corresponding NoDSs, and the last column depicts the reproduced scenarios validating that the original optimal paths can be traversed in the NoDSs.
V-A2 Root Causes of NoDSs
To better understand the detected NoDSs, we manually summarize six patterns based on the potential root causes for these non-optimal decisions. Table II shows the six unique root causes:
  1. 1.
    The ADS inaccurately predicts a moving NPC vehicle as stationary, leading to re-plan a non-optimal path.
  2. 2.
    The ADS inaccurately enlarge the impact region of static obstacles on the ego vehicle’s lane boundaries, resulting in the change of the optimal path.
  3. 3.
    The ADS inaccurately estimates accessible regions between parallel obstacles that do not obstruct the optimal path.
  4. 4.
    The ADS overly prioritizes an unrelated NPC’s movements, reducing feasible space for optimal path decision.
  5. 5.
    The ADS overly focuses on keeping a safe distance from surrounding static obstacles, resulting in insufficient space to select the optimal path.
  6. 6.
    The ADS is overly cautious about trailing vehicles, misclassifying part of the optimal area as non-optimal, resulting in insufficient space for executing the optimal path.
TABLE II: Proportion of discovered NoDSs
No. Root Cause Prop.(%) Uni. Uni.
1 Inaccurate prediction of vehicle status 6
2 Inaccurate prediction of obstacle impact 29
3 Inaccurate estimation of accessible region 28
4 Inaccurate prediction of other vehicles’ intentions 15
5 Excessive attention to surrounding obstacles 10
6 Excessively cautious about nearby vehicles 12
Table II shows the distribution of NoDSs discovered by Decictor, with the respective proportions for the six root causes being 6%, 29%, 28%, 15%, 10%, and 12%, demonstrating Decictor’s capability to identify a diverse range of root causes in NoDSs. The Uni. column underscores that Decictor identifies five unique root causes (No. 1 to 5) compared to the seven existing safety-critical ADS testing baselines. Notably, our method detects two unique NoDSs (No. 1 and 4) not identified by the Random-based baseline, Random-$\delta$ and Random, thus highlighting Decictor’s enhanced ability to discover varied root causes in NoDSs.
V-A3 Case Study
Fig. 6 shows two representative NoDS examples. More NoDS examples and corresponding videos can be found on our website [40]. The examples illustrate that the ego vehicle takes non-optimal paths in the detected NoDSs, even though the optimal ones are still available.
@@ V-A3 Case Study
Case 1. Inaccurate Prediction of Vehicle Status. The Decictor adds a new NPC vehicle in the NoDS, initially slow and located at a distance to the ego vehicle. As shown in the third column of Fig. 6, by the time the ego vehicle approaches, the NPC has moved, yet the original optimal path is still available. However, due to an incorrect prediction treating the NPC vehicle as stationary, the ADS stops lane changing and chooses a less efficient path.
@@
Case 2. Inaccurate Prediction of Obstacle Impact. Decictor places two obstacles near the junction exit, not affecting the optimal path. However, the ego vehicle (i.e., 2.1-meter width) deems the target lane (i.e., 3.5-meter width) impassable, choosing a suboptimal path from an adjacent lane, which raises risks in real-world traffic scenarios.
Answer to RQ1: Decictor significantly outperforms existing methods in identifying NoDSs. The non-invasive mutation can accurately generate valid scenarios (96.6% - 99.4%).
V-B RQ2: Usefulness of Mutation and Feedback
@@
We evaluated the key components of Decictor, including the non-invasive mutation and feedback mechanism, by configuring a series of Decictor variants.
V-B1 Mutation
@@
For mutation, we compared Decictor with its three variants: (1) w/o Cons applies a random mutation instead of the non-invasive mutation in Decictor, aiming to evaluate the effectiveness of the non-invasive mutation; (2) w/o Mot sets the time step to 0s in the calculation of non-invasive feasible areas (i.e., only considering the motion constraint at each timestamp), aiming to measure the influence of the motion constraint between two successive timestamps. (3) w/o Rem uses only the adding operation in Decictor under the constraint of non-invasive mutation, aiming to assess the effectiveness of the combination of the two mutation operations, i.e., adding and removing participants. As shown in Table III, we find that w/o Cons generates the fewest valid mutations (69.2%) and detects the smallest number of NoDSs (25.8), underlying the importance of the non-invasive mutation. Comparing w/o Mot with Decictor, we observed that only considering the motion constraint at each timestamp is insufficient, and the continuous-time motion constraint plays a significant role in calculating the non-invasive feasible area. Our preliminary study revealed that %Mutation increases while the #NoDS decreases as the time step increases. The reason is that a long time step will yield scenarios with reduced interactivity between the ego vehicle and the added participant, thus exerting a lower impact on the PPD process of the ADS. Comparing w/o Rem and Decictor, we found that mutation with only the adding operation is more likely to induce more invalid mutations and a lower number of NoDSs. This is because the adding operation introduces too many obstacles, potentially leading to the failure of the motion task. Thus, the removing operation is also important for generating NoDSs.
TABLE III: Effectiveness of mutations
Metric Method S1 S2 S3 S4 S5 S6 Avg. Sum
%Mutation\,$\uparrow$ w/o Cons 71.7 71.3 74.8 65.5 67.0 64.7 69.2 -
w/o Mot 71.1 80.9 98.6 68.9 77.6 66.1 77.2 -
w/o Rem 97.5 97.3 99.4 97.2 91.8 94.0 96.2 -
Decictor 98.0 97.5 99.4 97.5 96.6 98.2 97.9 -
#NoDS\,$\uparrow$ w/o Cons 2.7 4.5 1.6 5.7 7.0 4.3 - 25.8
w/o Mot 4.0 8.6 5.2 10.8 11.7 5.0 - 45.3
w/o Rem 3.9 9.3 4.9 9.1 8.3 5.7 - 41.2
Decictor 9.4 11.9 9.0 15.7 12.1 5.8 - 63.9
TABLE IV: Comparison of feedback
Method #NoDS\,$\uparrow$
S1 S2 S3 S4 S5 S6 Sum
F-Random 3.0 5.1 4.4 7.2 10.5 5.2 35.4
F-Con 3.0 6.7 1.2 7.5 11.7 4.9 35.0
F-Path 4.3 10.8 5.0 10.6 12.1 5.4 48.2
F-Behavior 5.3 9.0 5.4 11.5 11.8 5.6 48.6
Decictor 9.4 11.9 9.0 15.7 12.1 5.8 63.9
V-B2 Feedback
@@
For feedback, we implemented four variants: (1) F-Random replaces the feedback of Decictor with a random selection (from $\textbf{Q}\cup\textbf{Q}^{\prime}$) to evaluate the effectiveness of our feedback strategy. (2) F-Con replaces the feedback of Decictor with the grid similarity used in our consistency checking (see Equation 6) to compare the performance between using a coarse-grained grid distance and a fine-grained path distance. (3) F-Path and (4) F-Behavior consider only the driving path feedback and the behavior feedback, respectively, to evaluate the usefulness of either feedback type.
@@
Table IV shows the experimental results of #NoDSs. The comparative results between F-Random and Decictor (35.4 vs. 63.9 in total) illustrate the effectiveness of the feedback used in Decictor. The results of F-Con (35.0) indicates that the grid similarity is not an effective feedback metric, as it may overlook some scenarios with a high probability of inducing NoDSs. From the ablation results of F-Path and F-Behavior (48.2 and 48.6, respectively), we see that both behavior feedback and driving path feedback are beneficial for detecting NoDSs. Their combination achieves the best performance.
Note that, we also evaluated the influence of different $\Delta t$ used in non-invasive mutation. Due to the space limit, the detailed experimental results can be found on our website [40].
@@
V-C RQ3: Test Efficiency of Decictor
Figure 7: #NoDSs over the execution of different methods.
TABLE V: Results of time performance (s)
@@ V-C RQ3: Test Efficiency of Decictor
Fig. 7 shows the cumulative number of NoDSs over the execution time of different methods. We can find that as Decictor continues its execution, the detection of NoDSs steadily increases, whereas other methods quickly reach a stable state. We further assess the time performance of different components in Decictor, including the overhead of mutation, oracle checking, feedback calculation, and simulation. Specifically, we analyze the average time to handle a scenario. The results are summarized in Table V, where the columns Mutation, Oracle Check, Feedback, and Simulation represent the average time taken by a scenario for mutation, validity checking, fitness computation, and running the scenario in the simulation platform, respectively. 0.01* represents that the time cost is small, i.e., less than 0.01. For Random, the oracle and feedback stages are not involved, being marked as ‘N/A’.
The overall results show that, for all tools, the simulation process spends the majority of time. For instance, on average, Decictor takes 77.58 seconds to process a scenario, where 67.20 seconds (86.62%) are used in running the scenario. Decictor spends slightly more time in Mutation and Oracle Check than others due to the computation of the non-invasive feasible area for each waypoint and the grid-based similarity-checking mechanism. Note that we only consider the mutation time when generating scenario configurations. However, the computation time of Decictor remains within an acceptable range. SAMOTA consumes the most time in the feedback due to the necessity of training a surrogate model. The feedback times for scenoRITA and DoppelTest are approximately twice as long as those reported in [15]. This is due to the simulation length was set to twice the original settings to ensure the ego could complete the task. The differing time duration for simulation across various tools can be attributed to the different scenarios each tool generates (e.g., time-out scenarios), which in turn require varying amounts of simulation time.
V-D Threats to Validity
@@
Decictor suffers from some threats. First, the selection of the seed ODSs could potentially influence the results. To mitigate this, we have judiciously selected six motion tasks and created corresponding scenarios. The decision optimality of each ODS was manually analyzed and confirmed by all authors, including a co-author with industrial expertise. Second, the threshold selection for the Consistency Check $\epsilon$ in Decictor presents another threat, as different thresholds can lead to varying interpretations of non-optimal path decisions. To counter this, on one hand, we carefully select the threshold $\varepsilon$ in the Consistency Check and $\Delta t$ in the computation of non-invasive feasible area (the empirical evaluation of the two parameters can be found on our website [40]) to reduce the number of such scenarios; on the other hand, to ensure robustness and consistency, all co-authors independently review the generated NoDSs, and only those confirmed by unanimous agreement among all authors will be considered in NoDS-Hum. Third, the non-determinism inherent in ADS execution may also pose a threat to our results, and we mitigate it by repeating each experiment multiple times.
Lastly, the environments used could pose a potential threat. On one hand, we adapted existing baselines to the Apollo+SimControl simulation environment, which could potentially influence the results. However, our modifications were mainly limited to the interface between the algorithms and the simulation environment, leaving the core algorithms unchanged. We thoroughly reviewed the code and released the implementation of baselines on our website [40].
VI Related Work
@@
Safety-Guided ADS Testing. Safety plays the most important role in the development of ADSs. However, guaranteeing everlasting safety is challenging and many approaches have been proposed to generate safety-critical scenarios for evaluating the safety of ADSs. They can be divided into data-driven approaches [3, 4, 44, 45, 46, 47, 6, 48] and searching-based approaches [9, 49, 7, 50, 8, 10, 11, 33, 36, 51, 52, 15, 53]. Data-driving methods produce critical scenarios from real-world data, such as traffic recordings [4, 6, 47, 54] and accident reports [44, 45, 46, 3, 55, 56]. Search-based methods use various technologies to search for safety-critical scenarios from the scenario space, such as guided fuzzing [9, 8, 57, 58], evolutionary algorithms [7, 50, 52, 51, 33, 36, 16], metamorphic testing [59], surrogate models [10, 11], reinforcement learning [43, 60, 17] and reachability analysis [49, 61]. All these methods aim to generate critical scenarios for evaluating ADSs’ safety-critical requirements, such as collision avoidance and reaching the destination. In contrast, our work evaluates the non-safety-critical requirements of ADSs, which is also crucial for the real-world deployment of ADSs. Compared to these works, the key difference is that Decictor is the first to evaluate non-safety-critical requirements for ADSs, particularly in the aspect of non-optimal path-planning decisions, whereas the other works mainly focus on safety-critical issues.
Robustness Analysis for ADS. Robustness is a crucial issue for autonomous driving systems (ADSs). Recent studies have shown that the perception module in ADSs (i.e., object detection) exhibit vulnerabilities to various threats including perturbing sensor signals [62, 63], modifying objects [64, 65, 66], etc. There are also some studies [67, 68] work on the robustness of the prediction module by crafting trajectories of other vehicles. However, no existing work has tested the path-planning robustness of autonomous driving systems. In this paper, we propose the first work to reveal non-optimal PPDs of the ADS, which can be meaningful to inspiring the improvement on ADS robustness and reliability.
Path-planning in ADSs. Obtaining optimal PPDs is a challenging task in ADSs [69]. Existing studies on path planning primarily focus on avoiding obstacles, overlooking other attributes that contribute to optimality (e.g., motion time, path length, and comfort). Traditional algorithms [70, 71, 72, 73] often result in discontinuous paths, leading to non-optimal overall decisions. While, machine learning-based planning algorithms [74, 75, 76] could become trapped in local minima, unable to find optimal solutions. Therefore, given an ADS, it is critical to evaluate the robustness of its optimal PPDs, but there is still a significant gap in this area. To mitigate this, we proposed the first PPD testing technique.
VII Conclusion